Green Landscaping Inc. is preparing its budget for the first quarter of 2020. The next step in the budgeting process is to prepare a cash receipts schedule and a cash payments schedule. To that end the following information has been collected. Clients usually pay 60% of their fee in the month that service is performed, 30% the month after, and 10% the second month after receiving service. Actual service revenue for 2019 and expected service revenues for 2020 are November 2019, $93,660; December 2019, $82,580; January 2020, $104,570; February 2020, $124,710; and March 2020, $133,440. Purchases of landscaping supplies (direct materials) are paid 60% in the month of purchase and 40% the following month. Actual purchases for 2019 and expected purchases for 2020 are December 2019, $17,570; January 2020, $16,320; February 2020, $17,320; and March 2020, $22,350.

v9XrbbrWw2m1XXdruVt7e3RZ4XAI+PwIcpscAHAN
Answer #1

Schedule of cash payment

For the quarter ended 2020

January February March Quarter
December 7028 7028
January 9792 6528 16320
February 10392 6928 17320
March 13410 13410
Total payment 16820 16920 20338 54078

1) Account receivable = (124710*10%+133440*40%) = 65847

2) Account payable = 22350*40% = 8940